Two angles are supplementary. One angles is 30 les than five times the other angle. What are the measures of the angles?
Answer by stanbon(75887)

Two angles are supplementary. One angle is 30 less than five times the other angle. What are the measures of the angles?
----------------------------
If a and b are supplementary, a+b=180
EQUATION:
1st: a = 5b-30
2nd a+b = 180
----------------
Substitute 1st into 2nd to get:
5b-30 + b = 180
6b = 210
b = 35 degrees
a = 180-35 = 145 degrees
=============================
Cheers,
Stan H.